News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 10 November 2000 7.30am EST

Last week the observing schedule was interrupted following a high radiation event at 6:26pm EST on Sept 8 that stopped the daily load and activated the science instrument safing sequence (SCS 107). All actions were nominal and assessment of the dump data confirmed that the trigger was due to high proton counts in the EPHIN P4GM channel exceeding the pre-defined threshold of 300 counts/cm^2/s/sr. Observations will be resumed once the radiation falls to an acceptable level and impacted observations will be rescheduled.

A routine check of the CPE Warm Start Count was completed on 6 Nov and showed the value remained at 1.

A milestone was passed on November 1 - the official transition from observing cycle 1 to cycle 2 targets. November 1 was also the cut-off date for Cycle 1 peer reviewed Targets Of Opportunity. A small number of Cycle 1 targets have been scheduled for December and January due to their observing constraints.

The planned observations for the coming week are as follows.

------------------------------------------
Radiation Belts                      Nov 13 
Z Andromedae            ACIS-S/HETG
Hydra A                 ACIS-S/HETG
MCG-5-23-16             ACIS-S/HETG
RX J0925.7-4758         ACIS-S/HETG  Nov 14 
Radiation Belt                       Nov 15 
CTI                     ACIS         Nov 16 
A383                    ACIS-I
MITG J0414+0534         ACIS-S
A383                    ACIS-S
HCG 16 DOS              ACIS-S
M31 Nucleus             ACIS-S
M31 (multiple)          HRC-I
AC114                   ACIS-S       Nov 17
HE2149-2745             ACIS-S
Radiation Belts                      Nov 18 
Eta Car                 ACIS-S/HETG
PMN J1131+0455          ACIS-S       Nov 19 
NGC 3368                ACIS-S       Nov 20

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports